The half-life of an isotope is 10.0 minutes. If 25.0 grams are left after 60.0 minutes, how many grams were in the original sample?



1
Expert's answer
2021-12-07T05:37:02-0500

The half-life of a radioactive isotope is 10.0 min, i.e, "t_{1\/2}" = 10.0 min.

Therefore, the radioactive decay constant is "λ = \frac{0.693}{t_{1/2}} = \frac{0.693}{10.0 \;min}"

Let N0 is the initial amount of the radio isotope and Nt = 25.0 g is the amount remaining after t = 60.0 min.

Use the radioactive decay law.

"N(t) = N_0 \times e^{-λt} \\

ln (\frac{N_t }{N_0}) = -λ \times t"

Plug in values and get

"ln (\frac{25.0 }{N_0}) = -\frac{0.693}{10.0 \; min} \times 60.0 \; min = -4.158 \\

\frac{25.0 }{N_0} = e^{-4.158} \\

\frac{25.0 }{N_0} = 0.015638 \\

N_0 = 1598.67 \; g"

The amount of the radioactive isotope originally present is 1598.67 g (ans).